how do you get the answer to this problem?
1
- + 5=10
x

Answer by Alan3354(69443)

how do you get the answer to this problem?
1
- + 5=10
x
-----------
(1/x) + 5 = 10
Subtract 5
1/x = 5
Multiply by x
1 = 5x
Divide by 5
x = 1/5